OpenStudy (anonymous):

express as a single logarithm with a coefficient of 1 1/2 logb x - logb (1+x^2)

OpenStudy (anonymous):

is it to the base b??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/2 logb x = logb (x)^1/2 and use the formula log a - log b = log(a/b)

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

welcome...